The velocity distribution for flow over a flat plate is given by u = 3/2 y - y^(3/2) , where u is the point velocity in metre per second at a distance y metre above the plate. Determine the shear stress at y = 9 cm. Assume dynamic viscosity as 8 poise.
1
Expert's answer
2020-11-12T18:26:30-0500

Given "u=\dfrac{3}{2}y-y^{3/2}." Then


"\dfrac{du}{dy}=\dfrac{3}{2}-\dfrac{3}{2}y^{1/2}"

At "y=9 cm=0.09m"


"\dfrac{du}{dy}=\dfrac{3}{2}-\dfrac{3}{2}(0.09)^{1/2}=1.05(s^{-1})"

Shear stress is given


"\tau=\mu\cdot\dfrac{du}{dy}"

Assume dynamic viscosity as "\mu=8\ \text{poise}=0.8\ N\cdot s/m^2"

Then


"\tau=0.8\ N\cdot s/m^2\cdot1.05s^{-1}=0.84 N/m^2"

The shear stress at y = 9 cm is "0.84N\/m^2."